OpenStudy (anonymous):

how do I dilute 0.5 moles of zinc sulfate into 0.1 moles if the volume of the solution needs to be constant at 100 cm3?

OpenStudy (anonymous):

alright i'm assuming you are talking about a molar solution which molar = moles/litters, and you should also know that 1cm^3 =1ml So we know that \[100cm^3 (\frac{1ml}{1cm^3})(\frac{.1 mole zn}{1000ml zn})(\frac{1000ml}{.5 mole zn})\] So know we do our factor lablel cancell units and end up with 20 ml So we know it will take 20 ml of the .5 molar solution to dilute and get 100ml of the .1 molar sol.

OpenStudy (jfraser):

A handy shortcut for diluting solutions is\[M{_1}V{_1} = M{_2}V{_2}\] where 1 is the concentrated starting solution, and 2 is the more dilute solution. In this case, M1 is 0.5M, M2 will be 0.1M, and V2 will be 0.1L. Rearranging to solve for V1 gets you\[V{_1} = V{_2}\left[ M{_2}/M{_1} \right]\]\[V{_1} = 0.1L \left[ 0.1M/0.5M \right] = 0.02L\] which is 20mL.